# Parameter uncertainties for imperfect surrogate models in the low-noise regime

### v3

@@ II Analysis of the generalization error in the near-deterministic limit
We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$. The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method Swinburne and Perez (2024)
${\mathcal{G}}_{0}[\pi]$ $=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}$ (12)
$=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
@@
Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$.
To express the final generalization error, we introduce a pointwise mass function
${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$
@@
II.1 POPS-constrained ensemble and hypercube ansatz
@@
These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$. In the supplementary materials Swinburne and Perez (2024) we explore a weighted ensemble ansatz
@@
The ensemble ansatz produces highly informative max/min bounds for any test point, providing an envelope for worst case errors, as we show in figure 1. However, even when employing an optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see Swinburne and Perez (2024)), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2. In addition, both resampling and storage are $\mathcal{O}(N)$, which can become problematic for the $N/P\to\infty$ limit of interest.
@@
To determine $\mathcal{H}(\mathcal{T}^{*}_{E})$ we use singular value decomposition on the POPS-ensemble $\mathcal{T}^{*}_{E}$ to obtain the right eigenmatrix ${\bf V}\in\mathbb{R}^{R\times P}$, $R\leq P$, ${\bf V}{\bf V}^{\top}=\mathbb{I}_{R}$, requiring $\mathcal{O}(P^{3})$ effort for evaluation. Uniformly sampling the full max/min hypercube of the projected ensemble $\tilde{\mathcal{T}}^{*}_{E}=\{{\bf V}{\bm{\Theta}}|{\bm{\Theta}}\in{\mathcal{T}^{*}_{E}}\}$ returns samples $\tilde{{\bm{\Theta}}}^{\top}{\bf V}$ of $\pi^{*}_{\mathcal{H}}$ for $\mathcal{O}(R)$ effort and $\mathcal{O}(RP)$ storage. We find this efficient hypercube posterior to give excellent bounding and prediction of test errors for $P\in[3,1600]$, as shown in figures 2 and 3. An additional advantage is the simplicity of implementation for linear models; an example Python code is provided in the SM Swinburne and Perez (2024).
II.2 POPS-constrained loss minimization for linear models
@@
${\bm{\Theta}}^{*}_{\mathcal{L}}$ $={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (19)
${\bf A}$ $=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}}.$
The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme Swinburne and Perez (2024) that will be developed further in future work. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible. For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (20) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors, which can then be used to construct the ensemble or hypercube ansatz given above. A simple python implementation is provided in the SM Swinburne and Perez (2024).
III Numerical Experiments
III.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
@@
We applied Bayesian ridge regression and our POPS-ansatz to regress a quadratic polynomial against a sinusoidal function, the constant, linear and quadratic features giving $P=3$ model parameters. The SM Swinburne and Perez (2024) shows the same procedure for a model with $P=6$ independent features. For these simple models it is possible to approximately minimize the generalization error $\mathcal{G}_{E}$ of some discrete model ensemble, providing we regularize with a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}=\sigma{\bm{\Sigma}}^{*}_{\mathcal{L}}$. With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$, whilst in the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable. We found the same small $\sigma$ instability when applying an ensemble approach from Masegosa (2020) which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large. These instabilities show existing misspecification-aware methods require careful hyperparameter tuning and low model dimension to avoid vanishing gradients or other numerical issues. Our POPS-ansatz, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
@@
Figure 1 shows results for $\sigma=1,1/3,1/6$, at the lower limit for which minimization converges. As expected from section I.4, Bayesian ridge regression significantly underestimates test errors even at the 99.7% confidence level (three standard deviations) which as discussed above will worsen with increasing $N/P$. The max/min range of the numerical ensemble shows similar behavior for $\sigma=1,1/3$ whilst at $\sigma=1/6$ test errors are bounded but significantly overestimated, which worsens as $\sigma$ decreases (whilst stable). This was also seen on higher dimensional datasets Swinburne and Perez (2024). Importantly, in all cases our POPS ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$ give excellent bounding of test errors and are stable in the near-deterministic limit $\sigma\to 0$. As expected, the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ gave slightly more conservative bounds.
@@
Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points. Top left: mean and $3\sigma$ interval from Bayesian ridge regression. All other plots show mean and max/min range. Top: POPS-ensemble and POPS-hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$. Bottom: numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=1,1/3,1/6$. Lower $\sigma$ values gave numerical instabilities.
III.2 Test errors in medium-dimensional linear regression
@@
We now consider higher/medium dimensional problems, $P=10-50$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible. The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered Swinburne and Perez (2024).
As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound. Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$. The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios. We conclude with an application of the hypercube ansatz to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
@@
III.3 High-dimensional models for atomic simulations
Figure 3: Application to interatomic potentials. Left: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
Atomic-scale simulations of materials were traditionally broadly classified into two types:
@@ III.3 High-dimensional models for atomic simulations
The advent of ML has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019). The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range. Training an ML model then proceeds by minimizing a squared error loss between reference energies and gradients. While certain ML approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications. Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t. a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models. This regime corresponds to the underparametrized deterministic setting considered here. Estimating the errors incurred by such ML models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) .
In the following, a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ was used to predict per-atom contributions. Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 20, up to a maximal value of $N=129,853$. The properties of the ensemble are then evaluated on a complementary set of 10,000 hold-out testing points. Error analysis is carried out in the context of the POPS-hypercube ansatz. Even on the full dataset, the analysis requires on the order of 20 minutes on 1 CPU core. It is therefore extremely lightweight in addition to being easily parallelizable, enabling its use for extremely large datasets.

### v5

@@ II Analysis of the generalization error in the near-deterministic limit
We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$. The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method (appendix C)
${\mathcal{G}}_{0}[\pi]=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$ (12)
@@
Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, i.e. that the POPS is not empty, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, i.e. that no one parameter choice can be a member of all POPS. To express the final generalization error, we introduce a pointwise mass function
${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$ (14)
@@
III POPS-constrained ensemble and hypercube ansatz
@@
These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$. In appendix D we explore a weighted ensemble ansatz
@@
The ensemble ansatz produces highly informative max/min bounds for any test point, providing an envelope for worst case errors, as we show in figure 1. However, even when employing the variational optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see appendix D.1), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2. In addition, both resampling and storage are $\mathcal{O}(N)$, which can become problematic for the $N/P\to\infty$ limit of interest. As a result, whilst further investigation of ensemble schemes and alternatives to the variational optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (appendix D.1), we now look for an alternative ansatz beyond the ensemble approach.
@@
To determine $\mathcal{H}(\mathcal{T}^{*}_{E})$ we use singular value decomposition on the POPS-ensemble $\mathcal{T}^{*}_{E}$ to obtain the right eigenmatrix ${\bf V}\in\mathbb{R}^{R\times P}$, $R\leq P$, ${\bf V}{\bf V}^{\top}=\mathbb{I}_{R}$, requiring $\mathcal{O}(P^{3})$ effort for evaluation. Uniformly sampling the full max/min hypercube of the projected ensemble $\tilde{\mathcal{T}}^{*}_{E}=\{{\bf V}{\bm{\Theta}}|{\bm{\Theta}}\in{\mathcal{T}^{*}_{E}}\}$ returns samples $\tilde{{\bm{\Theta}}}^{\top}{\bf V}$ of $\pi^{*}_{\mathcal{H}}$ for $\mathcal{O}(R)$ effort and $\mathcal{O}(RP)$ storage. We find this efficient hypercube posterior to give excellent bounding and prediction of test errors for $P\in[3,1600]$, as shown in figures 2 and 3.
III.1 POPS-constrained loss minimization for linear models
@@
${\bf A}=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}},\quad{\bm{\Theta}}^{*}_{\mathcal{L}}={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (20)
The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme (appendix D.1) that will be developed further in future work. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible. For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of Ref. Mahalanobis, 1936. The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (21) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors, which can then be used to construct the ensemble or hypercube ansatz given above. Details of the implementation are provided in appendix E. In appendix F we compare the performance of our POPSRegression implementation against BayesianRidge from sklearn, finding an minimal overhead of around 2×2$\times$2 × for fitting and essentially identical performance for prediction, even when returning max/min bounds in addition to standard deviations.
IV Numerical Experiments
In this section we apply our POPS approach to a variety of synthetic and real datasets, comparing our POPS-ensemble and POPS-hypercube ansatz to Bayesian inference and the minimization of the PAC-Bayes generalization error $\mathcal{G}_{E}$. We also discuss our approach in the context of existing methods for uncertainty quantification in atomistic machine learning and discuss how these parameter uncertainties can be propagated up scales to predict material properties with uncertaintyMaliyov et al. (2024).
IV.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
@@
We applied Bayesian ridge regression and our POPS-ansatz to regress a quadratic polynomial against a sinusoidal function, the constant, linear and quadratic features giving $P=3$ model parameters. Appendix G shows the same procedure for a model with $P=6$ independent features. For these simple models it is possible to approximately minimize the generalization error $\mathcal{G}_{E}$ of some discrete model ensemble, providing we regularize with a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}=\sigma{\bm{\Sigma}}^{*}_{\mathcal{L}}$. With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$. In the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable, meaning that the final minimization is highly sensitive to initial conditions and frequently diverges, leading to conditioning issues. We found the same small $\sigma$ instability when applying an ensemble approach from Ref. Masegosa, 2020 which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large. These instabilities show existing misspecification-aware methods require careful hyperparameter tuning and low model dimension to avoid vanishing gradients or other numerical issues. Our POPS-ansatz, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
@@
Figure 1 shows results for $\sigma=1,1/3,1/6$, at the lower limit for which minimization converges. As expected from section I.4, Bayesian ridge regression significantly underestimates test errors even at the 99.7% confidence level (three standard deviations) which as discussed above will worsen with increasing $N/P$. The max/min range of the numerical ensemble shows similar behavior for $\sigma=1,1/3$ whilst at $\sigma=1/6$ test errors are bounded but significantly overestimated, which worsens as $\sigma$ decreases (whilst stable). Importantly, in all cases our POPS ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$ give excellent bounding of test errors and are stable in the near-deterministic limit $\sigma\to 0$. As expected, the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ gave slightly more conservative bounds. In appendix G we provide a similar test of directly minimizing the generalization error for higher dimensional datasets, with $P=6$ independent features. As for the polynomial case, we find that strong regularization is required for minimization to be stable, and in this limit the solution closely follows that of loss minimization. At weaker regularization, the direct minimization becomes systematically closer to the prediction of the POPS ansatz $\pi^{*}_{\mathcal{H}}$. These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
@@
Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points. Top left: mean and $3\sigma$ interval from Bayesian ridge regression. All other plots show mean and max/min range. Top: POPS-ensemble and POPS-hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$. Bottom: numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=1,1/3,1/6$. Lower $\sigma$ values gave numerical instabilities and are not presented (see appendix D).
IV.2 Test errors in medium-dimensional linear regression
@@
We now consider higher/medium dimensional problems, $P=10-100$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible. The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in a cubic polynomial. We also investigated the influence of adding a simple Gaussian noise term to simulates the presence of variability orthogonal to the chosen feature set. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation. In appendix G we provide examples showing this performance remains when varying the form of simulation engine to quadratic and sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models.
@@
As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound. Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$. The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios.
IV.3 Application to machine learning interatomic potentials
We now demonstrate the performance of the method on challenging high-dimensional $P=\mathcal{O}(100-1000)$ regression tasks from atomic machine learning. Atomic-scale simulations of materials were traditionally broadly classified into two types:
  1. a)
@@ IV.3 Application to machine learning interatomic potentials
The use of machine learning techniques in atomic simulation has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019). The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range. Training a machine learning model proceeds by minimizing a squared error loss between reference energies and gradients. While certain approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications. Further, quantum calculations can be tightly converged, so that their intrinsic error (w.r.t. a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models. This regime corresponds to the underparametrized deterministic setting considered here. Estimating the errors incurred by such machine learning models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021). In the following, we explore the application of our POPS approach for a linear interatomic potential model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$, with forces obtained by taking spatial gradients. Whilst we choose a particlar form of linear interatomic potential model, the presented approach is general to the wide range of available implementationsLysogorskiy et al. (2021); Wood and Thompson (2018); Podryabinkin and Shapeev (2017); Goryaeva et al. (2021).
Figure 3: Fitting a qSNAPWood and Thompson (2018) interatomic potential to a diverse tungsten datasetKarabin and Perez (2020). Left: representative training configurationsMontes de Oca Zapiain et al. (2022). Center: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
Our first application is to pure tungsten, a key material for nuclear fusion applicationsGoryaeva et al. (2021). Atomic training data was generated using a recently introduced information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022) that aims to maximize an estimated feature entropy for a given set of feature functions, giving an an extremely (ideally, maximally) diverse training set. 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components). Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), known as qSNAP, which can (in spite of what the name suggests) be expressed in linear model form $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ by defining the ${\bf F}_{i}$ as products of standard SNAP features. Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 21, up to a maximal value of $N=129,853$.